A barrel of crude oil has a volume of 42 gallons, only approximately 45% of which is processed into gasoline. If your car achieves 33 mi/gal, and you drive 36,000 miles in one year, how many barrels of crude oil are required to run your car for a year?


1
Expert's answer
2021-04-14T06:14:53-0400

Volume of a pure gasoline produced from one barrel of crude oil equals "42gal\times0.45=18.9gal"

Therefore,

"\frac{36000mi}{1year}\times\frac{1gal}{33mi}\times\frac{1barrel}{18.9gal}=58barrels/year"


Answer: 58 barrels